OpenStudy (anonymous):

Please Help! William is remodeling is home. In order to save money, he is upgrading the plumbing his self. He needs 24 pieces of cooper tubing, each 5/6 of a foot long. If he has a 17 foot piece of tubing will that be enough? how much more does he need or how much tubing will he have left?

OpenStudy (anonymous):

@study100 do you think you could help me out with this one?

OpenStudy (anonymous):

No he doesnt have enough tubing and he will need 3 1/3 more

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

First you need to find how many feet in total are the 24 pieces of copper 24* (5/6) = 20 ft. Now that he needs 20 ft, that means he needs: 20-17= 3 ft more

OpenStudy (anonymous):

3 feet or 3 1/3

OpenStudy (anonymous):

24* 5 = 120 ft 120/6 = 20 ft.

OpenStudy (anonymous):

Answer is 3 feet.

OpenStudy (anonymous):

Your right! thank you a bunch! i appreciate your help :)))

OpenStudy (anonymous):

No problem, glad it helps! <3